The longer side of a parallelogram is 54 c.m and the corresponding altitude is 16 c.m .if the altitude corresponding to the shorter side is 24 c.m  find the length of the shorter side
 

Grade:6

1 Answers

RUSHIKESH RAMESHCHANDRA BENKE
2 years ago
let, x be the length of the shorter side
we know that area of parallelogram=base * height
so longer side area and smaller side area are equal
therefore we comparing both are.we get
                          54 * 16=24 * x 
                     (54*16)24=x
                                    x=36
Length of smaller side is 36